OpenStudy (anonymous):

Can p^3 - 3p^2 + 5p - 15 be factored?

OpenStudy (anonymous):

my guess is no.

OpenStudy (raden):

put x=3, check it would be the zero :)

OpenStudy (anonymous):

but you can't factor anything correct?

OpenStudy (raden):

i u put x=3, gives 3^3-3(3)^2+5(3)-15=27-27+15-15=0 it means one of the factor of that polynomial is (x-3) so, this can be factored :)

OpenStudy (raden):

* if

OpenStudy (anonymous):

how did you figure 3?

OpenStudy (raden):

i have used by trials, first :)

OpenStudy (raden):

another way, u can use wolfram to check it http://www.wolframalpha.com/input/?i=x^3-3x^2%2B5x-15